name three verbs used in doing mathematics with examples


1
Expert's answer
2022-02-14T16:21:01-0500

"To simplify" means to reduce an expression into fewer terms or into a form that is easier to work with.

For example, 2x = 4 can be simplified into x = 2.

"To evaluate" an algebraic expression means to find the values of the expression when the variable is replaced with a given number.

For example, we should evaluate 2x − 4 when x = 3.

"To factor", or factoring, means to break an expression down into its factors.

For example, 2x + 16 = 2×(x + 8).